Fermi National Laboratory


Accelerator Update

Friday October 17 through Monday October 20
All shifts began in shutdown.

On Friday after noon, two Linac experts accessed their beamline and discovered a small LCW leak on their Lambertson (BLAM). This leak is internal to the magnet. The Technical Division said they would repair the leak on Monday morning. On Friday evening Operations received an alarm for the Linac Buncher. Operators investigated and found that Linac line voltage took a dip at the same time. Unfortunately, the Buncher now has a problem. Technicians will look at it on Monday. The evening shift had been having I- Source trips all evening. Operators called an expert who expained a procedure for conditioning. Operations had the source working normally by the end of the shift.

On Saturday, early on the evening shift, two Booster RF stations (BRF13 & 14) tripped off twice. The both reset okay each time. Then BRF16 tripped twice and later BRF11 tripped off. They all reset okay.

On Sunday, early on the midnight-shift a power supply for a Klystron station (KRF4) tripped and wouldn't rest. Then a Linac RF station (LRF3) tripped off. It also wouldn't reset. An expert said to leave it off and techs would fix it on Monday. During the evening shift BRF13 &14 tripped off due to a vacuum burst. It reset okay.

On Monday, during the midnight shift, operators found several LCW leaks in the Antiproton Source tunnel. The Main Injector had a planned ten-minute power outage at 7 AM this morning for maintenance work.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/20/2003   email Fermilab